OpenStudy (anonymous):

Integral of 1/t^2(t-1) dt

OpenStudy (anonymous):

Is the answer \[\left(\begin{matrix}1 \\ t\end{matrix}\right) - \ln \left(\begin{matrix}t \\ t-1\end{matrix}\right) + c\] the same as \[\frac{ -2 }{ t^3(t-1) } + \frac{ \ln(t-1) }{ t^2 } + c\]

OpenStudy (anonymous):

I used t^2 = u and dv = 1/(t-1)

OpenStudy (thomas5267):

What is the integral? \(\int\frac{1}{t^2(t-1)}dt\)? What is \(\binom{1}{t}\) and \(\ln\binom{t}{t-1}\)? Binomial coefficients?

OpenStudy (anonymous):

the correct answer is \[\frac{ 1 }{ t } - \ln \frac{ t }{ t-1 }\] and what I have written as the last equation is my solution

OpenStudy (anonymous):

and yes thats teh integral

hartnn (hartnn):

you did integration by parts ?

hartnn (hartnn):

when both the functions of of same type, here algebraic, integration by parts is not recommended. for this problem, partial fractions is the easiest way. know what partial fractions is ??

OpenStudy (anonymous):

yes I made u = t^-2 and dv = (t-1)^.-1= 1(/(t-1)

OpenStudy (thomas5267):

I have no idea how to tackle this integral. \(t=\sec^2(x)\)?

OpenStudy (anonymous):

I don't by terms probably, but i've probably done it, could u explain?

hartnn (hartnn):

sure \(\huge \dfrac{1}{t^2(t-1)}=\dfrac{A}{t}+\dfrac{B}{t^2}+\dfrac{C}{t-1}\) you just need to find A,B,C

OpenStudy (anonymous):

Oh yes!, I have done this, okey let me try to see if I can solve this.

hartnn (hartnn):

take your time :)

OpenStudy (anonymous):

okey so A = 2 B and C = -1?

hartnn (hartnn):

doesn't seem to be correct, how did u get them ?

hartnn (hartnn):

1 = At (t-1) +B (t-1) +C (t^2) plug in t=0, you will get B plug in t= 1, you will get C

OpenStudy (anonymous):

-b = 1 that gives b = -1. Sorry first of the equation 1 = A(t()t-1) + b(t-1) + c(t^2) and now after getting b, I have 1 = -At -Bt and if t = 1 u have 1 = -A-B so -A = 1+B hat gives A = -1-B= -1-(-1) = 0 so I was wrong here okey, and C, 1 = At^2+ Ct^2 and if t=1 then 1 = A+C which gives C equals to 1, I was wrong here again. Conclusion, A = 0, B = -1, C = 1

hartnn (hartnn):

how can u say -b=1 ?

OpenStudy (anonymous):

B has a term that doesnt include t B(t-1) which is Bt and -B, but you way of doing this was easier.

hartnn (hartnn):

oh yes!, B=-1 is correct :) find A and C now

hartnn (hartnn):

C= 1 is also corect

hartnn (hartnn):

but A is Not 0

OpenStudy (anonymous):

A =-1

OpenStudy (anonymous):

when t = 2

OpenStudy (anonymous):

But the answer in my book is \[\frac{ 1 }{ t } - \ln \frac{ t }{ t-1 } + c\]

hartnn (hartnn):

you will get that :) plug in A,B,C values!

OpenStudy (anonymous):

\[\frac{ -1 }{ t }- \frac{ 1 }{ t^2 }\ + \frac{ 1 }{ t-1 }\]

OpenStudy (anonymous):

and now?

hartnn (hartnn):

yes now integrate!

hartnn (hartnn):

-1/t^2 is -t^-2 its integral will be ?

OpenStudy (anonymous):

oh okey, 1 sec

OpenStudy (anonymous):

-t^-1 = -1/(t)

hartnn (hartnn):

thats integral of just t^(-2)

OpenStudy (anonymous):

okey hold on

OpenStudy (anonymous):

t^-1 = 1/(t)

hartnn (hartnn):

yes, correct and what about other 2 terms, you can integrate those ?

OpenStudy (anonymous):

yes, give me a sec:)

OpenStudy (anonymous):

okey so for -t^-1 the answer is -ln t for 1/(t-1) its ln (t-1)?

hartnn (hartnn):

correct

hartnn (hartnn):

factor out the negative -ln t + ln (t-1) = - [ ln t - ln (t-1)]

hartnn (hartnn):

= - ln [t/(t-1)] :)

OpenStudy (thomas5267):

God... I was thinking how to use trig substitution for \(\int\frac{1}{t^2}dt\)...

OpenStudy (anonymous):

oh right thats how you do it, I gotta tell you ever since I have progressed in the math courses I've started to felt more stupid, is that normal?

OpenStudy (anonymous):

feeö*

OpenStudy (anonymous):

feel

hartnn (hartnn):

ofcourse...you're learning, silly mistakes are bound to happen, and thats a good thing! :)

OpenStudy (anonymous):

Well, huge thanks to both of you and for taking ur time ;)

hartnn (hartnn):

welcome ^_^ always happy to help :)

OpenStudy (anonymous):

hehe, well you will probably see more of me later or sooner, so I'll go ahead and close this thread now:)